### v1

# A Statistician’s Overview of Mechanistic-Informed Modeling

Abstract
The recent success of deep neural network models with physical constraints (so-called, Physics-Informed Neural Networks, PINNs) has led to renewed interest in the incorporation of mechanistic information in predictive models. Statisticians and others have long been interested in this problem, which has led to several practical and innovative solutions dating back decades. In this overview, we focus on the problem of data-driven prediction and inference of dynamic spatio-temporal processes that include mechanistic information, such as would be available from partial differential equations, with a strong focus on the quantification of uncertainty associated with data, process, and parameters. We give a brief review of several paradigms and focus our attention on Bayesian implementations given they naturally accommodate uncertainty quantification. We then specify a general Bayesian hierarchical modeling framework for spatio-temporal data that can accommodate these mechanistic-informed process approaches. The advantage of this framework is its flexibility and generalizability. We illustrate the methodology via a simulation study in which a nonlinear Burgers’ equation PDE is embedded via a Bayesian PINN.
Keywords Bayesian, Burgers, Deep Learning, Hierarchical, Spatio-Temporal, Uncertainty Quantification
1 Introduction
As Galileo supposedly noted, “Mathematics is the language in which God has written the universe.” Indeed, much of our understanding of the world around us, including physical laws and principles, is codified in terms of mechanistic formulas, such as partial and ordinary differential equations. From a statistical perspective, one can attempt to infer mechanistic equations that describe the system under study. Moreover, one can use a priori knowledge of mechanistic equations describing a physical process of interest, along with data, for inference (e.g., learning the parameters that control the specified equations or initial conditions) or prediction (e.g., forecasting/interpolating).
When data are involved, one inherently has uncertainty when building models and optimization methods become an integral part of understanding the world around us. The classic example is when Carl Friedrich Gauss used the method of least squares in the early 1800s to successfully calculate the orbits of celestial bodies; importantly, Gauss used Kepler’s laws of celestial mechanics as constraints to his least squares solution (Gauss,, 1809). Indeed, statisticians have long been interested in using proposed mathematical relationships in the presence of observations to motivate statistical models. For example, as described in Wikle & Hooten, (2010), in 1927, Yule used pendulum motion differential equations to suggest an autoregressive time series for sunspot observations and Hotelling used approximations of differential equations to develop models for population growth in the United States. The integration of mechanistic information in statistical models is now a fundamental component of several sub-disciplines of statistics; e.g., statistical meteorology (Berliner et al.,, 1999), ecological statistics (Hefley et al.,, 2017), epidemiology (Leach et al.,, 2020), glaciology (Gopalan et al.,, 2018), statistical oceanography (Herbei et al.,, 2008; Milliff et al.,, 2011), pharmacokinetic and pharmacodynamic analysis in the development of pharmaceuticals (e.g., Gabrielsson & Weiner,, 2001), to name a few.
With the advent of deep learning, the use of mechanistic information combined with deep neural network models has seen an exponential increase in applications across many disciplines. Such approaches, often known as Physics Informed Neural Networks (PINNs, Raissi et al., 2017a, ), are typically based on augmenting the neural network loss (e.g., mean squared error) with loss terms corresponding to the physical equations and boundary/initial conditions (e.g., Karniadakis et al.,, 2021). Although this is novel in the context of neural network models, augmenting loss functions to include mechanistic information when fitting models to data is not new. For example, Sasaki pioneered the approach in a variational solution to the data assimilation problem in meteorology beginning in the 1950s (e.g., Sasaki,, 1958, 1970). Around the same time, control engineers were combining (linearized) dynamical equations within a filter that updated in the presence of new observations (e.g., Kalman,, 1960). Additional approaches developed in engineering and statistics to “fuse” data, random processes, and deterministic models were developed starting in the 1990s and continue to this day (e.g., see overviews in Berliner,, 2003; Cockayne et al.,, 2019; Meng et al.,, 2020).
From a statistical perspective, the goal is not to develop a “solver” for a known mechanistic set of equations as is often of interest in the PINN literature (Raissi et al.,, 2019), but rather to recognize, in the G. Box sense, that all our mechanistic models are in some way “wrong” when applied to real world scenarios, and that we need to account for the uncertainty in the model specification, controlling parameters, and data. In this sense, the Bayesian hierarchical modeling (BHM) approach has long been seen as the natural framework to account for all of these sources of uncertainty in physics-motivated statistical modeling (e.g., see Berliner et al.,, 1999; Cressie & Wikle,, 2011; Hennig et al.,, 2015). The Bayesian view has also increasingly been shared in the PINN community (e.g., Yang et al.,, 2021; Meng et al.,, 2022; Huang et al.,, 2023; Psaros et al.,, 2023). This has been an important development from a statistical perspective. Yet, some components of the BHM modeling paradigm are often not explicitly represented in these implementations, and there is still very little consideration of the importance of specifying dependent random effects, which has been shown to be very useful in statistical mechanistic modeling of spatio-temporal processes (e.g., Wikle et al.,, 2001; Cressie & Wikle,, 2011; Chkrebtii et al.,, 2016).
Our goal in this paper is to present a statistically-minded overview of mechanistically-informed modeling in the presence of data and uncertainty. Note, we prefer the term “mechanistically-informed” to “physics-informed” because many mechanistic models for non-physical processes (e.g., invasive species in ecology, Wikle, (2003)) also follow this paradigm and it can be misleading to use the term “physics” for such processes. We will approach this from a general BHM perspective, incorporating both neural and statistical surrogate models for the mechanistic process (in our case, PDEs). We first provide a brief historical overview of how mechanistic process information has been combined with data in various modeling paradigms. Throughout, we discuss issues of interest to statisticians, such as uncertainty in data, parameters, and process knowledge, as well as computational tractability. We then show how the PINN perspective fits naturally into a general BHM framework.
Section 2 discusses the general problem at hand and introduces the relevant notation. Then, in Section 3 we present a brief review of mechanistically-informed modeling from the perspective of variational data assimilation and optimal interpolation, as well as statistical approaches, with a focus on Bayesian and hierarchical Bayesian implementations to account for uncertainty. Section 4 describes PINNs and Bayesian-PINNs (B-PINNs) as well as connections to the aforementioned earlier literature. Section 5 discusses a general BHM framework that can accommodate mechanistically-driven process models (e.g., physical-statistical models and B-PINNs). We then demonstrate in Section 6 this general approach with a B-PINN corresponding to a Burgers’ equation mechanistic PDE via a simulation study. We conclude with a brief overview and discussion of challenges in Section 7.
2 Problem Outline
Our interest in mechanistically-informed models here concerns problems common to spatio-temporal statistics where one has potentially incomplete and noisy observations of a univariate or multivariate process at some locations in space and time, and seeks to: (a) determine an initial condition (i.e., the true process values at specified spatial locations at some initial time), (b) predict (interpolate) incomplete observations at specified locations in space and time, or (c) forecast at specified spatial locations at times beyond which we have observations. We generally assume that we have some a priori knowledge of the underlying spatio-temporal dynamics of the process, yet our knowledge is incomplete. That is, we may not know the controlling parameters of the mechanistic model and/or the mechanistic model is an incomplete representation of the process. (An incomplete representation of dynamics could be due to many circumstances – for instance, due to approximations to the underlying partial differential equations, or in other instances due to missing mechanistic understanding.) Next, we outline the notation to be used in this exposition.
2.1 Process and Data Notation
Assume our interest is in a $K$-variate process at spatial locations $\bm{s}\in\mathcal{D}$ (in $\mathbb{R}^{1},\mathbb{R}^{2}$, or $\mathbb{R}^{3}$) and times $t\in[0,T+\tau]$, denoted by $\bm{u}_{t}(\bm{s})\equiv(u^{(1)}_{t}(\bm{s}),\ldots,u^{(K)}_{t}(\bm{s}))^{\prime}$. Ultimately, we are interested in this process at some fixed set of $n$ spatial locations, say $\mathcal{S}\equiv\{\bm{s}_{1},\ldots,\bm{s}_{n}\}$ and some discrete set of $n_{\mathcal{T}}$ times $\mathcal{T}\subseteq\{0,1,\ldots,T+\tau\}$ (although, the approach presented here can accommodate continuous time). If we stack these vectors, we can write the $Kn$-vector $\bm{u}_{t}\equiv(\bm{u}_{t}(\bm{s}_{1})^{\prime},\ldots,\bm{u}_{t}(\bm{s}_{n})^{\prime})^{\prime}$. For example, if the process is univariate then $K=1$ and $\bm{u}_{t}\equiv(u_{t}(\bm{s}_{1}),\ldots,u_{t}(\bm{s}_{n}))^{\prime}$. More generally, observations could be available at different locations for each of the $K$ variables of interest, but we do not add that additional notational complexity here.
Assume we have observations of elements of the process corresponding to some or all of the times in $\mathcal{T}$ at $m_{t}$ spatial locations $\bm{r}_{j}\in\mathcal{D}_{m}\subseteq\mathcal{D}$, which we denote by the $m_{t}$-vector $\bm{z}_{t}$. Some or all of these observation locations may coincide with the spatial locations in $\mathcal{S}$, but may not, as discussed below. Note, it may be the case that $m_{t}=0$ for some times (i.e., there are no observations for certain times).
2.2 Mechanistic Model Notation
Here we focus our attention on mechanistic models that take the form of a time-dependent partial differential equation (PDE), but the ideas presented here can be easily adapted to other forms of mechanistic models (e.g., ordinary differential equations, or agent-based models). Assume we have a PDE that we believe is mechanistically informative to our latent process, $\bm{u}_{t}(\bm{s})$. For example, we may believe that our latent process shows evidence of advection and diffusive dynamics. We denote this
$\frac{\partial\tilde{\bm{u}}_{t}(\bm{s})}{\partial{t}}=\mathcal{M}_{u}(\tilde{\bm{u}}_{t}(\bm{s}),\bm{\theta}_{m}),$ (1)
where $\tilde{\bm{u}}_{t}(\bm{s})$ corresponds to the $K$-dimensional mechanistic process defined at $t\in[0,T+\tau]$ and $\bm{s}\in\mathcal{D}$ and $\mathcal{M}_{u}$ is an operator controlled by parameters $\bm{\theta}_{m}$. Note, we are deliberately making a distinction between the true latent process of interest, $\bm{u}_{t}(\bm{s})$ (no tilde), and the mechanistic model representation, given by $\tilde{\bm{u}}_{t}(\bm{s})$ this distinction will be made more clear when we discuss models. Assume that the PDE has associated initial and boundary conditions, respectively,
$\tilde{\bm{u}}_{0}(\bm{s})$ $=\bm{g}_{I}(\bm{s}),\;\;\bm{s}\in\Omega,$
$\bm{B}(\tilde{\bm{u}}_{t}(\bm{s}))$ $=0,\;\;t\in[0,T+\tau],\bm{s}\in\mathcal{D}_{B},$
where $\bm{g}_{I}(\cdot)$ is some function defined at the initial time, 0, $\bm{B}(\cdot)$ is a function defined on the boundary $\mathcal{D}_{B}$.
Our goal is to perform inference on $\bm{u}_{t}(\bm{s})$ at the $\mathcal{S}$ spatial locations and at times $t\in\mathcal{T}$ while taking into account the information present in $\mathcal{M}_{u}$ and given the data $\bm{Z}\equiv\{\bm{z}_{1},\ldots,\bm{z}_{T}\}$. We can denote this as a posterior distribution (with some abuse of notation) as $[\bm{u}_{t}:t\in\mathcal{T}|\bm{Z},\mathcal{M}_{u}]$, where we use the bracket notation ([][\;\;][ ]) to represent a probability distribution, as is common in the Bayesian hierarchical modeling literature (e.g., Cressie & Wikle,, 2011); that is, $[a|b]$ corresponds to the distribution of $a$ given (conditional upon) $b$, $[b]$ corresponds to the marginal distribution of $b$, and $[a,b]$ corresponds to the joint distribution of $a$ and $b$.
In practice, we may be interested in learning the initial condition ($\bm{u}_{0}$), interpolation at some time $t$ ($\bm{u}_{t}$), or forecasting ($\bm{u}_{T+\tau}$). This is generally termed the “forward problem” where $\mathcal{D},\mathcal{D}_{B},\bm{g}_{I}(\cdot),\bm{B}(\cdot)$, and $\bm{\theta}_{m}$ are known and one seeks to find $\bm{u}_{0}(\bm{s})$ and $[\bm{u}_{t}]_{t\in[0,T+\tau]}$ given the dynamics suggested by $\mathcal{M}_{u}$. We may also be interested in learning parameters associated with the mechanistic PDE ($\bm{\theta}_{m}$) and various parameters associated with the underlying statistical or neural models used to represent the PDE (see below). This is generally termed the “inverse problem,” where we may have partial observations of the process in the domain of interest. In practice, we are often interested in learning about both $[\bm{u}_{t}]_{t\in[0,T+\tau]}$ on $\mathcal{D}$ and $\bm{u}_{0}(\bm{s})$ on $\mathcal{D}_{B}$ as well as the $\bm{\theta}_{m}$ that satisfies $\mathcal{M}_{u}$. This is a major focus of the methodology presented below.
In the exposition that follows, we will sometimes refer to the numerical solution of the PDE in the short-hand notation,
$\tilde{\bm{u}}_{t+1}=\mathcal{M}(\tilde{\bm{u}}_{t};\bm{\theta}_{m}),$ (2)
but recognize that there are PDE solvers that provide $\tilde{\bm{u}}_{t}$ for any and all times in $[0,T]$ that are not necessarily sequential in time (e.g., implicit solvers). Note, we drop the subscript $u$ on $\mathcal{M}$ to distinguish this numerical operator from the theoretical operator given in (1).
3 Historical Perspective
Fundamental advances in mechanistic-informed modeling in the presence of data originated with the operational need to assimilate observations with physical models of the atmosphere for weather forecasting (i.e., numerical solutions of simplified Navier-Stokes PDEs). The main purpose of such “data assimilation” is to generate initial conditions for the numerical forecast models that are consistent with the dynamics of the system, but these initial conditions have incomplete observations in space (e.g., we cannot observe the atmosphere everywhere). We discuss the data assimilation approach in Section 3.1. The connection to more statistical approaches and further developments are discussed in Section 3.2.
3.1 Variational Data Assimilation Approach
It was known shortly after L.F. Richardson’s failed attempt at the first numerical weather forecast in the early 1920s (Richardson,, 1922) that the numerical solution of PDEs with initial conditions given by data would quickly become unstable due to lack of consistency with the governing equations. Thus, by the time that computers were actually capable of generating a realistic forecast in real time, it was known that initial condition fields of weather variables had to be smoothed in a way that respected dynamical constraints (Charney et al.,, 1950; Phillips,, 1956). Two main approaches were developed in the 1950s and 1960s to do this, one deterministic based on the calculus of variations and one stochastic, based on what would come to be known as Gaussian processes (GPs). These approaches are essentially equivalent when one takes a Bayesian inferential perspective (e.g., Wikle & Berliner,, 2007). In this section, we briefly discuss these approaches as they provide the foundation for mechanistically-informed models for spatio-temporal processes in statistics.
The initial idea in data assimilation that one could simultaneously minimize an objective function containing both the sum of squared differences between the observations and the “true” initial field as well as a constraint that incorporated the PDE dynamics was presented in Sasaki, (1958). Although he formulated this in continuous space, practical uses of the methodology considered discrete space (e.g., Thompson,, 1969; Sasaki,, 1970; Stephens,, 1970).
Following the notation in Section 2.1, say we have $m_{t}$-dimensional data vectors, $\bm{z}_{t}$, and numerical PDE solutions of the latent (state) process at time $t$ denoted by $\tilde{\bm{u}}_{t}$ given by (2). Here, without loss of generality, we assume $K=1$ (a univariate process) for notational simplicity. Critically, for deterministic systems, if one knows the initial state $\tilde{\bm{u}}_{0}$ then (2) determines the sequence of future states, $\{\tilde{\bm{u}}_{1},\tilde{\bm{u}}_{2},\ldots,\tilde{\bm{u}}_{t},\ldots\}$. In Sasaki’s variational approach, one then breaks the objective function into two components $J_{0}(\tilde{\bm{u}}_{0})$ and $J_{1}(\tilde{\bm{u}}_{0})$, given as follows (e.g., see Lewis & Lakshmivarahan,, 2008). First, let
$J_{0}(\tilde{\bm{u}}_{0})=\frac{1}{2}\sum_{t=1}^{T}(\bm{z}_{t}-\bm{h}(\tilde{\bm{u}}_{t}))^{\prime}\bm{R}_{t}^{-1}(\bm{z}_{t}-\bm{h}(\tilde{\bm{u}}_{t})),$ (3)
where $\bm{h}(\cdot)$ is a (known) mapping function that takes the $n$-dimensional deterministic model state vector $\tilde{\bm{u}}_{t}$ and maps to the $m_{t}$-dimensional observation vector at time $t$ (e.g., an interpolator if data are not coincident with model locations, or perhaps the identity if they are; this function can also take into account nonlinear transformations, see Daley, (1993)). In addition, $\bm{R}_{t}$ is the time-varying $m_{t}\times m_{t}$ observation error covariance matrix (representing measurement error and other sources of observational error), also typically assumed known. The physical constraint is then given by:
$J_{1}(\bm{\lambda},\tilde{\bm{u}}_{0})=\sum_{t=1}^{T}\bm{\lambda}_{t}^{\prime}(\tilde{\bm{u}}_{t+1}-\mathcal{M}(\tilde{\bm{u}}_{t})),$ (4)
where $\bm{\lambda}_{t}$ are $n\times 1$ vectors of Lagrange multipliers and $\bm{\lambda}\equiv\{\bm{\lambda}_{1},\ldots,\bm{\lambda}_{T}\}$. Thus, the optimal initial state vector is found by minimizing the loss function: $L(\bm{\lambda},\tilde{\bm{u}}_{0})=J_{0}(\tilde{\bm{u}}_{0})+J_{1}(\bm{\lambda},\tilde{\bm{u}}_{0})$, where we note that $J_{1}$ is implicitly a function of the initial state, $\tilde{\bm{u}}_{0}$, through the dynamic model. Indeed, the constraints here come through the dynamic model, $\mathcal{M}(\cdot)$. Given that this model is deterministic and assumed to be “perfect”, this is considered a “strong” constraint –i.e., the solution must follow the deterministic model (Sasaki,, 1970).
3.1.1 Operational Variational Data Assimilation (4DVar)
In practice, variational data assimilation evolved away from Lagrange multipliers to an unconstrained approach known as “4DVar” (“4D” because it includes 3 dimensions in space and time). As with the Sasaki approach, we have data in the interval $[1,T]$ and seek the initial condition, $\tilde{\bm{u}}_{0}$. The observation component of the objective function is the same as (3) above where again the $\tilde{\bm{u}}_{t}$ are functions of the initial condition $\tilde{\bm{u}}_{0}$ and obtained deterministically from the numerical solution of the PDE (2), which is assumed to be error free. In this approach, one then assumes that there is some background “estimate” of the initial condition, say $\bm{u}_{b}$, such as a forecast from the previous assimilation cycle, or some historical average. Then, the background term of the objective function is given by,
$J_{1b}(\tilde{\bm{u}}_{0})=\frac{1}{2}(\tilde{\bm{u}}_{0}-\bm{u}_{b})^{\prime}\bm{C}_{b_{0}}^{-1}(\tilde{\bm{u}}_{0}-\bm{u}_{b}),$ (5)
where $\bm{C}_{b_{0}}$ is the background error covariance matrix (a spatial covariance matrix for the errors between the true initial state and the previous forecast of the initial state). The optimization minimizes the sum of $J_{0}+J_{1b}$ (i.e., equation (3) plus (5)) . Importantly, including the dynamical constraints implicitly through the deterministic model integration effectively converts the constrained minimization problem (as in (4)) to an unconstrained minimization problem. In practice, the optimization is facilitated by using an adjoint method (e.g., Lewis & Derber,, 1985; Errico,, 1997).
3.1.2 Weak Constraint 4DVar
One can consider weak-constraint variational assimilation by incorporating an explicit model error term in the objective function (e.g. Bennett,, 1992, via the method of representors). For example, say we have the model error term, $\bm{\eta}_{t}$, as in
$\bm{u}_{t+1}=\mathcal{M}(\bm{u}_{t};\bm{\theta}_{m})+\bm{\eta}_{t},$ (6)
where $\bm{\eta}_{t}\sim\text{Gau}(\bm{\mu}_{\eta},\bm{Q})$, where the mean $\bm{\mu}_{\eta}$ vector and error covariance matrix $\bm{Q}$ are assume known or estimated offline. In this section, we prefer $\bm{u}_{t}$ instead of $\bm{\tilde{u}}_{t}$ since $\bm{u}_{t}$ is not constrained to follow the numerical model exactly, as in the previous subsections. Then, the extra term in the objective function in the 4DVar case is given by (e.g. Laloyaux et al.,, 2020),
$J_{2}(\bm{u}_{0})=\frac{1}{2}\sum_{t=1}^{T}(\bm{\eta}_{t}-\bm{\mu}_{\eta})^{\prime}\bm{Q}^{-1}(\bm{\eta}_{t}-\bm{\mu}_{\eta}),$ (7)
where $\bm{\eta}_{t}\equiv\bm{u}_{t+1}-\mathcal{M}(\bm{u}_{t};\bm{\theta}_{m})$ from (6) and again, $\bm{u}_{0}$ is implicitly included in each of the model error terms, $\bm{\eta}_{t}$. Minimizing the loss function $J_{0}+J_{1b}+J_{2}$ (i.e., equation (3) plus (5) plus (7)) in terms of $\bm{u}_{0}$ is constrained by the numerical model but can deviate from an exact solution of the model, resulting in a weak constraint.
3.1.3 Optimal Interpolation
Around the same time that the variational approach to meteorological data assimilation was being developed, other researchers were developing statistical optimization approaches based on stochastic processes. A brief history of these approaches is presented in Lewis & Lakshmivarahan, (2008). Although there is some debate about its origins, Lewis & Lakshmivarahan, (2008) describe the first practical approach to come from A. Eliassen in an obscure technical report from Oslo University in 1954. More formal development was given by Gandin, (1963). Gandin’s approach is widely considered to have been equivalent to the formalization of kriging by Matheron (Matheron,, 1963), which is the essence of geostatistical modeling (“kriging”) in spatial statistics; see Cressie, (1990) for a discussion of the origins of kriging. In retrospect, all these approaches simply use Gaussian processes to perform interpolation (spatial prediction) that is optimal in the sense of minimizing the expected squared error loss (i.e., risk).
In the context of meteorological data assimilation, one interpolates the residuals by removing a background field (e.g., the previous forecast) from observations. So, given an observation vector, $\bm{z}_{t}$, and a forecast from the previous time, $\bm{u}_{b}$ (where again, $b$ stands for “background”; note, keeping with the notation used in this paper, when this forecast is from the numerical solution of a PDE, we could denote the background as $\tilde{\bm{u}}_{b}$), the Eliassen (simple kriging) approach gives the latent state vector ${\bm{u}}_{t}$ from:
${\bm{u}}_{t}=\bm{u}_{b}+\bm{K}_{t}(\bm{z}_{t}-\bm{H}_{t}\bm{u}_{b}),$
where $\bm{H}_{t}$ is an $m_{t}\times n$ matrix that maps the state process to the observations at time $t$, and $\bm{K}_{t}$ is the $n\times m_{t}$ matrix of interpolation weights given by
$\bm{K}_{t}=\bm{C}_{b}\bm{H}_{t}^{\prime}(\bm{H}_{t}\bm{C}_{b}\bm{H}_{t}^{\prime}+\bm{R}_{t})^{-1},$
where $\bm{C}_{b}$ is the $n\times n$ spatial covariance matrix of the background (forecast) error. This matrix must be specified in terms of a covariance function, and because one may need to develop predictions from locations that are not observed, it comes from a covariance function associated with a Gaussian process (e.g., see Cressie & Wikle,, 2011).
There are different ways that physical constraints are included in this framework. First, note that the forecast for time $t$, which we are denoting by $\bm{u}_{b}$ here, comes from a forecast given by the dynamic model (2) and the previous estimate of the state, say ${\bm{u}}_{t-1}$. Alternatively, one can specify covariance functions that satisfy dynamical constraints, which is especially useful in multivariate applications of this methodology (e.g., see Daley,, 1993).
Note, this formulation does not explicity account for multiple observations over times; i.e., there is no dependence specified across time. In addition, this optimal interpolation approach for spatial fields can also be formulated from a variational perspective as in Section 3.1.1, known as “3DVar” in the data assimilation literature. Spatial optimal interpolation and 3DVar are equivalent if the observation mapping function is linear, i.e., $\bm{h}(\bm{u}_{t})=\bm{H}_{t}\bm{u}_{t}$.
3.1.4 Connections to Filtering and Bayesian Estimation
About the same time that Sasaki was developing the variational approach to combining data and mechanistic model information, control engineers were working on the same problem, from an online filtering perspective. The classic example of this is the Kalman filter (Kalman,, 1960), which has become ubiquitous in control engineering and statistical state-space time-series and spatio-temporal analysis (e.g., see Shumway & Stoffer,, 2000; Wikle & Berliner,, 2007). Typically in the control engineering case, one assumes they have an initial condition and initial variance/covariance matrix and proceeds to generate forecasts and update those as new observations come online, then generates another forecast, etc. The filtering equations explicitly account for model error, as in the weak constraint 4DVar case. Indeed, there is a clear connection between the filtering formula and the variational equations (e.g., see Wikle & Berliner,, 2007, for a derivation and connections).
As discussed in Wikle & Berliner, (2007), one can view the variational, optimal interpolation, and filtering approaches as Bayesian inference, where there is a likelihood model relating observations to a latent dynamic process, and then the process is assigned a prior distribution with the background vector (e.g., forecast) $\bm{u}_{b}$ as the prior mean. When the data and prior distributions are Gaussian and the mapping function is linear, the variational and Bayesian approaches are equivalent (e.g., Wikle & Berliner,, 2007).
One of the limitations of this perspective is that it assumes the model parameters (and error variances/covariances) are known (or they are somehow estimated off-line). In the case of online filters, one can augment the state vector with parameters and update them along with the state process. Alternatively, in variational approaches it is certainly possible to add another term to the objective function that accounts for parameter uncertainty. This is equivalent to assigning a prior distribution to the parameters. When practical, it is often more flexible to specify prior distributions explicitly in case non-Gaussian priors are warranted. The development of mechanistically-motivated spatio-temporal dynamic models (described in the next section) evolved from this hierarchical Bayesian perspective.
3.2 Statistical Approaches for Adding Mechanistic Information
There have been several approaches to combine mechanistic information in statistical spatial and spatio-temporal models. For example, the development of spatial and spatio-temporal covariance functions from analytical or numerical solutions to PDEs has long been considered an efficient way of incorporating mechanistic information in spatial and spatio-temporal models. PDE-derived covariance functions have been used for spatial prediction, such as for kriging models (e.g., Heine,, 1955; Whittle,, 1954), and have recently become quite popular with the stochastic-PDE (SPDE) approach (e.g., Lindgren et al.,, 2011, 2022). Indeed, as mentioned in Section 3.1.3, this idea of using PDE-derived covariance functions for optimal interpolation has also been used in the early data assimilation literature (e.g., Daley,, 1993). Generally, this approach is useful for developing covariance functions, but is quite limited in the complexity of the underlying mechanistic model that can be effectively considered (see the overview in Cressie & Wikle,, 2011).
As an alternative, in the functional data and non-parametric regression analysis literature, it has long been common to use function derivatives to regularize a regression model such as in smoothing splines for spatial data (e.g., Wahba,, 1990; Eubank,, 1999; Ramsay,, 2002). Over the last 10 years or so, this has been extended to the physics-informed regularization case in spatial and spatio-temporal regression (e.g., see the review in Sangalli,, 2021). These methods are very much in the spirit of the Sasaki, (1958) approach described in Section 3.1, where one adds a PDE regularization term to the loss function that compares the non-parametric model of interest to the data. As described in Tomasetto et al., (2024), point estimates of parameters can be estimated in this framework. These approaches can be quite effective for certain problems and in fundamental ways, the PINN approach described below in Section 4.1 can be thought of as a generalization of this approach with more flexible semi-parametric (deep neural) models and the use of automatic differentiation to allow more complex PDE regularization. Bayesian-PINNs (described in Section 4.2), mechanistically-motivated statistical models (described in Section 3.2.1), and probabilistic numerical methods (described in Section 3.2.2) operate similarly, but can accommodate more principled uncertainty quantification across all phases of model development via a hierarchical Bayesian framework. Given uncertainty quantification is a major focus for modern spatio-temporal modeling applications, we describe the statistical approaches in more detail below in the following subsections, and PINNs and B-PINNs in Section 4.
3.2.1 Mechanistically-Motivated (Physical-Statistical) Models
Before introducing the mechanistically-motivated framework, we first briefly review Bayesian hierarchical models for spatio-temporal processes. Adding additional sources of data and accounting for uncertainty in observation model parameters and state-process model parameters is often needed when modeling complex spatio-temporal data. In some cases, we can gain much more expressiveness in our models if we add dependence structure (time, space, covariates) to the parameters as well (i.e., treat them as dependent random processes, conditionally). Bayesian hierarchical models are natural extensions of the Bayesian paradigm that allow for more complete accommodation of uncertainty and additional model flexibility.
The Bayesian hierarchical (deep) modeling paradigm for spatio-temporal data (e.g., Berliner,, 1996; Cressie & Wikle,, 2011) considers breaking the joint distribution of data, process, and parameters into a series of conditional models (distributions). In that context, the posterior distribution is then given by the following decomposition
$[process,\;parameters\;|\;data]$ $\propto$ $[data|process,data\;parameters]$
×$\times\;\;\;[process|process\;parameters]$
×$\times\;\;\;[parameters].$
Importantly, each of these stages can have sub-stages as described below in Section 3.2.1.
In the context of the dynamic spatio-temporal models (DSTMs) of interest here, a basic BHM might take the following form.
Data Model: $\;\;\;\bm{z}_{t}=\bm{h}_{t}(\bm{u}_{t})+\bm{\epsilon}_{t},\;\;\bm{\epsilon}_{t}\overset{ind.}{\sim}\;\text{Gau}(\bm{0},\bm{R}_{t}(\bm{\theta}_{z})),$
Process (Dynamic) Model: $\;\;\;\bm{u}_{t}=\mathcal{M}(\bm{u}_{t-1},\bm{B}(\bm{u}_{t};\bm{\theta}_{b});\bm{\theta}_{m})+\bm{\eta}_{t},\;\;\bm{\eta}_{t}\overset{ind.}{\sim}\;\text{Gau}(\bm{0},\bm{Q}_{t}(\bm{\theta}_{q})),$
Process Initial Value: $\;\;\;\bm{u}_{0}\sim\text{Gau}(\bm{a},\bm{\Sigma}_{0})$
Parameter Models: $\;\;\;[\bm{\theta}_{z},\bm{\theta}_{m},\bm{a},\bm{\Sigma}_{0},\bm{\theta}_{b},\bm{\theta}_{q}],$
where $\bm{B}(\bm{u}_{t};\bm{\theta}_{b})$ represents a boundary condition and $\bm{u}_{0}$ is the initial condition, both of which depend on parameters that are given prior distributions in general. Importantly, the parameter models can also have dependence (in space and/or time) or depend on other covariates, which gives this model a great amount of generalizability and can make it much “deeper” in terms of having more levels of conditioning.
Including Mechanistic Information: The BHM framework naturally allows one to include mechanistic information in each model layer. Here, we briefly describe how mechanistic information can be incorporated into the data, process, and parameter levels.
Data: A strength of the BHM approach is that one can easily specify multiple data sources conditioned on the latent state process, e.g., $[Z_{1},Z_{2}|u]=[Z_{1}|u][Z_{2}|u]$ (assuming conditional independence). In statistics, using mechanistic model output (e.g., weather reanalysis fields) as “data” (say, $Z_{2}$) has been a common “data fusion” approach to add mechanistic information since the late 1990s (e.g., Berliner et al.,, 1999; Wikle et al.,, 2001; Berliner,, 2003; Fuentes & Raftery,, 2005; Wikle & Berliner,, 2005). This is a type of weak constraint mechanistic modeling that has the advantage of being able to account for additive errors (to facilitate uncertainty quantification).
Process: As with the data assimilation literature, one can specify $\mathcal{M}(\cdot)$ to be a known mechanistic model. Alternatively, in the physical-statistical (mechanistically-motivated) DSTM paradigm, one can specify statistical models, $\bm{M}(\cdot;\bm{\theta}_{m})$, that are “motivated” by mechanistic processes, but are conditioned on parameters ($\bm{\theta}_{m}$) that typically have spatial or temporal structure (e.g. spatially-varying diffusion parameters); see Wikle & Hooten, (2010) and Cressie & Wikle, (2011) for numerous examples. For example, one may discretize (or write in spectral/finite-element space) a reaction-diffusion PDE or integro-difference equation and condition this on the diffusion/growth parameters that may vary with space as a GP, or in terms of covariates, which are specified at the next level of the hierarchy (e.g., Wikle,, 2003; Xu et al.,, 2005; Hooten & Wikle,, 2008). If desired, one can include hard constraints (i.e., “mass balance”) in the physical-statistical framework (e.g. Berliner et al.,, 2008; Zhuang & Cressie,, 2014; Gopalan et al.,, 2018).
Parameters: In some cases, mechanistic information can be imparted to the parameters (say, $\bm{\theta}_{m}$) through the prior specification; e.g., regularization or informed by mechanistic information. For example, Wikle, (2003) allowed covariates and a latent Gaussian process to model the spatially-varying diffusion rate in reaction-diffusion model for spread of an invasive species and Wikle et al., (2001) included a known turbulent scaling relationship as a prior on the variances of tropical surface winds to add realistic structure. Finally, Leeds et al., (2014) used pre-training/meta-learning of prior distributions with model output obtained from a coupled ocean-ecosystem model to inform a quadratic nonlinear mechanistically-motivated statistical model for data assimilation of ocean color measurements.
3.2.2 Probabilistic Numerical Methods
Over the last decade, there has been growing interest in modernizing the ideas associated with treating differential equation solutions as random processes (e.g., Larkin,, 1972; Diaconis,, 1988). These ideas, which come under a broad set of methods called probabilistic numerical methods (PNMs), have provided a comprehensive toolkit to consider uncertainty quantification in the solution of differential equations (e.g., see Hennig et al.,, 2015; Chkrebtii et al.,, 2016; Cockayne et al.,, 2019, and references therein). These methods have the advantage of being quite rigorous in their convergence guarantees and are straightforward to implement.
Many PNM methods for solving PDEs with uncertainty assume the solution has a GP prior. Then observations of boundary conditions and forcing terms of the PDE at a finite set of points act as data that allow for a posterior distribution over the PDE solution. For a simple illustration, in the space-only case, we might have the following GP prior:
$u_{t}(s)\sim GP(m(s),k(s,s^{\prime};\theta_{k})),$
where $m(s)$ is a mean function and $k(s,s^{\prime};\theta_{k})$ is a covariance function between two spatial locations $s$ and $s^{\prime}$, defined by parameters $\theta_{k}$. Thus, this prior accommodates belief about the inherent smoothness of the PDE solution. In the case of a linear map (say, $\mathcal{M}(u(s))=f(s)$), the implied GP prior for the PDE solution is
$\mathcal{M}(u(\bm{s}))\sim GP(\mathcal{M}(m(s)),\mathcal{M}{\mathcal{M}}^{\prime}(k(s,s^{\prime};\theta_{k}))),$
where $\mathcal{M}{\mathcal{M}}^{\prime}(k(s,s^{\prime};\theta_{k}))$ is the quadratic form associated with the linear operation on the prior covariance function (i.e., the kernel derived based on the fact that $\bm{A}\bm{\Sigma}\bm{A}^{T}$ is the variance-covariance matrix of random vector $\bm{A}\bm{x}$ where the variance-covariance matrix of $\bm{x}$ is $\bm{\Sigma}$). One then treats forcing terms as data, e.g.,
$f(s_{i})=\mathcal{M}(u(s_{i}))+\epsilon_{i},\;\;\epsilon_{i}\sim\text{Gau}(0,\sigma^{2}_{\epsilon}),$
where $\sigma^{2}_{\epsilon}$ controls the fidelity of the PDE solution. Given the prior and data models are Gaussian, the posterior distribution of $u(s)$ given $f(s_{1}),\ldots,f(s_{m})$ are available in closed form via a Gaussian distribution. In practice, one includes a boundary condition process and, for time-varying PDEs, an initial condition. In addition, one typically chooses “collocation” points for which to obtain “observations” $f(s_{i})$ to facilitate computation. This approach is often embedded in a BHM with real world observations to learn the parameters of the underlying PDE (i.e., the inverse problem).
The strength of the PNM approach is its principled uncertainty quantification, which comes from the numerical approximation (e.g., the $f(s_{i})$ observations) and, in the case of inverse problems, real-world observations. The inherent regularization of the Bayesian prior also provides some robustness to ill-posedness. Perhaps most importantly, PNM’s strong connections to traditional numerical methods have been established for special cases (e.g., linear PDEs and GP priors). That said, these methods are much more difficult to implement for nonlinear PDEs, do not scale well for very high-dimensional problems, and have not been used as much in deep BHMs with multiple sources of process uncertainty, as in the physical-statistical paradigm. However, the Bayesian PINN framework described in the next section can be considered a neural network based PNM.
4 Physics-Informed Neural Networks (PINNs)
In the spirit of the Sasaki’s variational approach described above in Section 3.1, consider fitting a deep neural network but constraining the solution to follow a PDE. This is the essence of PINNs (Raissi et al., 2017a, ; Raissi et al., 2017b, ; Raissi,, 2018; Raissi et al.,, 2019, 2020). Such approaches can be used in the context of fitting a PDE in the presence of data (as in data assimilation), or solving a PDE explicitly (presumably for the sake of efficiency). We are interested here in the situation where we have data corresponding to the latent process of interest, initial conditions, or boundary conditions. We provide a brief overview of the PINN approach and its Bayesian counterpart (B-PINN) in this section, and then describe how these can fit into a mechanistically-informed BHM framework in Section 5.
4.1 PINN Formulation
PINNs, first proposed by Raissi et al., 2017a ; Raissi et al., 2017b , mainly consist of two components: (1) a neural network for prediction and (2) a physically-inspired equation (e.g., PDE) to impose constraints in the loss function. We denote the deep neural network solution to the PDE at time $t$ and location $\bm{s}$ as $\bm{u}_{NN,t}(\bm{s};\bm{\theta}_{W})$, where $\bm{\theta}_{W}$ represents the neural network parameters (weights and biases). Note, $\bm{u}_{NN}$ is an approximation of the PDE solution $\tilde{\bm{u}}$. Indeed, since this is not a perfect representation it is best to make the distinction as follows by defining a PDE “residual” (as in the PNM approach described in Section 3.2.2). Let $\bm{\theta}_{m}$ correspond to PDE parameters as in (1), and write
$\bm{r}_{t}(\bm{s};\bm{\theta}_{W},\bm{\theta}_{m})\equiv\frac{\partial{\bm{u}_{NN}(\bm{\theta}_{W})}}{\partial{t}}-\mathcal{M}_{u}({\bm{u}_{NN}(\bm{\theta}_{W})};\bm{\theta}_{m}),$ (8)
for the “residuals” of the PDE evaluated at time $t$ and location $\bm{s}$ (note, in some cases where there is an exogenous forcing term in the PDE, this “residual” could be thought of as a representation of the forcing). In practice, as in the PNM approach, one considers $N_{c}$ so-called collocation points at which to evaluate the PDE, say $\{\bm{s}_{c}^{i},t_{c}^{i}\}_{i=1}^{N_{c}}$. Note, there can be different numbers of spatial locations at different times, so the $N_{c}$ collocation points include the complete set of space/time locations to train the neural net on the PDE (although we typically are interested in the solution at other locations in time and space). The loss function associated with the PDE is then given by the mean squared error (MSE),
$J_{P}(\bm{\theta}_{W})=\frac{1}{N_{c}}\sum_{i=1}^{N_{c}}\|\bm{r}_{t_{c}^{i}}(\bm{s}_{c}^{i};\bm{\theta}_{W})\|_{2}^{2},$ (9)
where $\|\;\|_{2}$ represents the $\ell_{2}$ norm and, for now, we are assuming the PDE parameters $\bm{\theta}_{m}$ are known, and so exclude them notationally. We typically would also include similar terms for the initial and boundary conditions, i.e., where there are $N_{IC}$ initial condition collocation points and $N_{BC}$ boundary condition collocation points. Denote these initial and boundary collocation points as $\{\bm{s}_{IC}^{i}\}_{i=1}^{N_{IC}}$ and $\{\bm{s}_{BC}^{i},t_{BC}^{i}\}_{i=1}^{N_{IC}}$, respectively; then the corresponding loss functions can be written as:
$J_{IC}(\bm{\theta}_{W})=\frac{1}{N_{IC}}\sum_{i=1}^{N_{IC}}\|{\bm{u}}_{NN,0}(\bm{s}_{IC}^{i};\bm{\theta}_{W})-\bm{g}_{I}(\bm{s}_{IC}^{i})\|_{2}^{2},$
$J_{BC}(\bm{\theta}_{W})=\frac{1}{N_{BC}}\sum_{i=1}^{N_{BC}}\|\bm{B}({\bm{u}}_{NN,t_{BC}^{i}}(\bm{s}_{BC}^{i};\bm{\theta}_{W}))\|_{2}^{2}.$
If one is simply interested in finding the solution of the PDE given the initial and boundary conditions and PDE parameters $\bm{\theta}_{m}$, then one finds the $\bm{\theta}_{W}$ that minimize the combined objective function,
$J(\bm{\theta}_{W})=a_{IC}J_{IC}(\bm{\theta}_{W})+a_{BC}J_{BC}(\bm{\theta}_{W})+a_{P}J_{P}(\bm{\theta}_{W}),$
where the specified non-negative constants $\{a_{IC},a_{BC},a_{P}\}$ allow one to focus more on different components of the solution if desired. The solution in this case does not require measurements of the process besides initial and boundary condition training data, so the PINN serves as an alternative to the traditional numerical solver (e.g., finite difference, finite element or spectral scheme). In this context, note a well-known limitation of the PINN approach is it can have difficulty in simulating multi-scale, chaotic, or turbulent behavior because the explicit forward (“causal”) nature of time is not considered (e.g., Wang et al.,, 2022). This limitation can be mediated in a number of ways, for example, by adding a time weighting function in the loss function (9). Regardless, some recent comparison studies have shown that traditional finite element solutions of PDEs work better and are as or more efficient than PINNs for most problems (e.g., Grossmann et al.,, 2024).
4.1.1 PINNs with Observational Data
As mentioned previously, in most cases in statistical applications of mechanistic-informed models with neural networks, our interest is to fit a deep neural network model with PDE constraints to data, which in the PINN literature is considered to be “data-driven discovery” of the PDE, by which they mean to learn the PDE parameters, $\bm{\theta}_{m}$ (i.e., an inverse problem). In this case, a data component is added to the objective function as in the variational approach and the PDE parameters can be estimated jointly with neural network parameters by minimizing the associated loss; the loss function $J_{P}$ in (9) is a function of $\bm{\theta_{m}}$ in addition to $\bm{\theta_{W}}$ due to the dependence of $\bm{r}_{t}(\cdot)$ on $\bm{\theta_{m}}$. Moreover, additional data observations of the process can be used; i.e., given observations, $\{\bm{z}_{t^{i}}(\bm{s}^{i}):i=1,\ldots,N\}$,
$J_{D}(\bm{\theta}_{W})=\frac{1}{N}\sum_{i=1}^{N}\|\bm{z}_{t^{i}}(\bm{s}^{i})-\bm{u}_{NN,t^{i}}(\bm{s}^{i};\bm{\theta}_{W})\|_{2}^{2}.$ (10)
Note, in the original representation of data-driven discovery, there is no measurement error associated with the observation in (10). The solution proceeds by optimizing the weighted combination
$J(\bm{\theta}_{W},\bm{\theta}_{m})=a_{IC}J_{IC}(\bm{\theta}_{W})+a_{BC}J_{BC}(\bm{\theta}_{W})+a_{P}J_{P}(\bm{\theta}_{W},\bm{\theta}_{m})+a_{D}J_{D}(\bm{\theta}_{W}),$
where we now explicitly note that $J_{P}$ depends on the PDE parameters $\bm{\theta}_{m}$. After this optimization, one has learned the PDE parameters and the associated neural network can be used to obtain an estimate of $\bm{u}_{t}(\bm{s})$ at desired locations in space and time.
Importantly, “data-driven discovery of PDEs” does not mean finding a functional form of PDEs in this case, but rather indicates estimating parameters within PDEs. The use of statistical models and deep neural networks to learn the functional form of mechanistic models is a growing area of research, but beyond the scope of this article. For recent overviews, see North et al., (2023), Sahimi, (2024), and North et al., (2025).
4.2 Bayesian PINNs
The PINN approach described above does not explicitly account for uncertainty in the data, model, or parameters. It is quite straightforward to account for measurement error as in the variational approach presented earlier; we can modify the data portion of the loss function (10) by assuming that the measurement errors are Gaussian. For example, we might assume that we have observation models of a univariate process (although this can be generalized to a multivariate process) $z_{t^{i}}(\bm{s}^{i})=u_{NN,t^{i}}(\bm{s}^{i};\bm{\theta}_{W})+\epsilon_{t^{i}}(\bm{s}^{i}),$ where $\epsilon_{t^{i}}(\bm{s}^{i})\sim\text{Gau}(0,\sigma^{2}_{u})$. Similarly, in principle we could have observations of $r_{t}(\bm{s})$ (the PDE residual term), boundary condition $B(\cdot)$ and initial conditions $g_{I}(\cdot)$ that also have independent Gaussian errors and associated error variances. All these distributions are conditioned on the solution of the PDE, but in the PINN context, that is completely determined given the NN and PDE model parameters, $\bm{\theta}_{W},\bm{\theta}_{m}$. Note, in principle, one could have dependent errors as in the variational and physical-statistical approach above.
It is then natural, following from the Bayesian NN literature (e.g., Neal,, 2012), to specify prior distributions on the NN weight and bias parameters, $\bm{\theta}_{W}$; known as a B-PINN (e.g., Yang et al.,, 2021). That is, assume we generically have conditional data distributions for the process, PDE residual, boundary and initial conditions given by $[D_{u}|\sigma^{2}_{u},\bm{\theta}_{W},\bm{\theta}_{m}]$, $[D_{r}|\sigma^{2}_{r},\bm{\theta}_{W},\bm{\theta}_{m}]$, $[D_{bc}|\sigma^{2}_{bc},\bm{\theta}_{W},\bm{\theta}_{m}]$ and $[D_{ic}|\sigma^{2}_{ic},\bm{\theta}_{W},\bm{\theta}_{m}]$, respectively. Each of these terms is implicitly conditioned on $u_{NN,t}$, but as mentioned above, since the NN is completely determined conditioned on its weights and biases, it is not necessary to write this explicitly. In addition, as in the case of PNMs, unless we have a forcing term in the PDE, we assume implicitly that “observations” of the residuals are equal to zero on average – that is, conditioned on the parameters, the relationship in (1) holds perfectly up to some arbitrary white noise (represented by $\sigma^{2}_{r}$). This has the effect of including a term equivalent to (9) in the un-normalized posterior representation (12) below. That is, we can either write this as a “data model” with known zero observations or as a regularization prior term. For the latter, we might write the prior distribution as:
$\frac{\partial{u_{NN}(\bm{\theta}_{W})}}{\partial{t}}\sim\text{Gau}(\mathcal{M}_{u}({u_{NN}(\bm{\theta}_{W})};\bm{\theta}_{m}),\sigma^{2}_{r}),$ (11)
which corresponds to a regularization term and discrepancy parameter, $\sigma^{2}_{r}$, that effectively specifies a priori belief in the representativeness of the PDE for this process.
Assuming conditional independence of the different data types, one can express the conditional data distribution as
$[D_{u},D_{r},D_{bc},D_{ic}|\bm{\theta}_{W},\bm{\theta}_{m},\bm{\theta}_{D}]=[D_{u}|\sigma^{2}_{u},\bm{\theta}_{W},\bm{\theta}_{m}][D_{r}|\sigma^{2}_{r},\bm{\theta}_{W},\bm{\theta}_{m}][D_{bc}|\sigma^{2}_{bc},\bm{\theta}_{W},\bm{\theta}_{m}][D_{ic}|\sigma^{2}_{ic},\bm{\theta}_{W},\bm{\theta}_{m}],$
where $\bm{\theta}_{D}\equiv(\sigma^{2}_{u},\sigma^{2}_{r},\sigma^{2}_{bc},\sigma^{2}_{ic})^{\prime}$ and assume we have prior distributions $[\bm{\theta}_{W}]$, $[\bm{\theta}_{m}]$, and $[\bm{\theta}_{D}]$. Then, Bayes’ rule gives
$[\bm{\theta}_{W},\bm{\theta}_{m},\bm{\theta}_{D}|D]\propto[D|\bm{\theta}_{W},\bm{\theta}_{m},\bm{\theta}_{D}][\bm{\theta}_{W}][\bm{\theta}_{m}][\bm{\theta}_{D}],$ (12)
where $D\equiv\{D_{u},D_{r},D_{bc},D_{ic}\}$. The intractable posterior (12) is typically sampled from using MCMC methods (e.g., Hamiltonian Monte Carlo, Langevin Dynamics), variational methods, or ensemble methods (see the comprehensive review in Psaros et al.,, 2023).
A significant challenge with Bayesian NNs (and, hence, B-PINNs) is the specification of reasonable prior distributions for the NN parameters $\bm{\theta}_{W}$. In general, such parameters are not identifiable, which can make MCMC sampling challenging without extra constraints (e.g., McDermott & Wikle,, 2019), the specification of informative priors (e.g., through empirical Bayes or meta-learning; e.g., Meng et al., (2022)), or special classes of priors (e.g., Meng et al.,, 2022; Sell & Singh,, 2023; Castillo & Egels,, 2024).
Note that the B-PINN framework presented here does not include an explicit term for model error, although $\sigma^{2}_{r}$ does account for minor discrepancies in the PDE. But, this does not respect the fact that the PDE dynamics are not usually sufficient to model the real world process. In other words, the real-world dynamical process is latent and is motivated by the PDE, but is typically more complicated. This is exactly the motivation for denoting a distinction between the true latent process, $u$, and the PDE solution $\tilde{u}$ in Section 2.2 and the NN solution $u_{NN}$ here. Such discrepancy terms are common in the computer model experiment literature (e.g., Gramacy,, 2020). In the context of B-PINNs, a discrepancy term can be added in the data model, e.g., $z_{t^{i}}(\bm{s}^{i})=u_{NN,t^{i}}(\bm{s}^{i};\bm{\theta}_{W})+\delta_{t^{i}}(\bm{s}^{i};\bm{\theta}_{\delta})+\epsilon_{t^{i}}(\bm{s}^{i})$, where $\delta_{t^{i}}(\bm{s}^{i};\bm{\theta}_{\delta})$ is a discrepancy term that depends on parameters $\bm{\theta}_{\delta}$. In the statistical surrogate and mechanistically-motivated statistics literature, this discrepancy is typically modeled as a GP, a functional (basis) expansion, or a machine learning surrogate. In the B-PINN context, it is natural to model the discrepancy as an additional (separate) neural network (e.g., Zou et al.,, 2024), with the discrepancy used to correct for the misspecification of the PDE when observing PDE residuals. In either case, one must be careful to ensure that there is sufficient prior information regarding the spatial (and possibly temporal) scale of the discrepancy process to ensure identifiability of the solution (see our example in Section 6 below for more discussion).
Although the B-PINN framework is an increasingly useful paradigm to account for uncertainty in PINN applications, there are still potential limitations in its lack of an explicit hierarchical formulation, lack of accommodation of dependent random effects, focus on Gaussian data distributions, and lack of more complicated measurement/sampling errors. In the next section, we show how the PINN fits into a more general deep mechanistically-informed BHM framework that can accommodate these concerns.
5 General Mechanistically-Informed Deep Dyamic Spatio-Temporal Model (GMID-DSTM)
Here, we present a general mechanistically-informed deep dynamic spatio-temporal model (GMID-DSTM) for statistical analysis of spatio-temporal data. The paradigm is flexible enough to include the examples presented previously (e.g., physical-statistical models, B-PINNs) as special cases. Importantly, this framework also can account for non-Gaussian data, model output as data, dependent random effects, and covariate effects. We describe the approach below following the spatio-temporal hierarchical modeling principles outlined in Cressie & Wikle, (2011).
5.1 Data Models
Assume we have up to $p$ different datasets at time $t$ given by $\{\bm{z}_{1t},\ldots,\bm{z}_{pt}\}$ that are all thought to be informative about our latent process of interest, $\bm{u}_{t}$. Here, we assume these datasets are conditionally independent given the latent spatio-temporal process. Specifically, $\bm{z}_{jt}|\bm{u}_{t},\bm{\theta}_{zj}\sim\;\;\mathcal{D}(h_{j}(\bm{u}_{t});\bm{\theta}_{zj}),$ where for $j=1,\ldots,p$, $\mathcal{D}(h_{j}(\bm{u}_{t});\bm{\theta}_{zj})$ is some data distribution (e.g., Gaussian, Poisson, log-normal, etc.) that is conditioned on the latent spatio-temporal process, $\bm{u}_{t}$, and parameters $\bm{\theta}_{zj}$, and where $h_{j}(\;)$ is a mapping function between the latent process and the data that can account for missing observations, nonlinear transformations, and change-of-support (spatial resolution) (for more background see Cressie & Wikle,, 2011).
The important aspect of these data models is they can accommodate multiple data sources (data fusion) of different types (e.g., count data, positive support data, Gaussian, e.g., Bradley, (2022)), different supports (spatial resolution and alignment, e.g., Wikle & Berliner, (2005)), measurement errors, and missing data. In addition, as mentioned previously, one can include output from a mechanistic model (e.g., an earth-system model) as “data”, which is another way to incorporate mechanistic information into the analysis (e.g., see Wikle et al.,, 2001; Wikle & Berliner,, 2005; Fuentes & Raftery,, 2005; Leeds et al.,, 2014, for examples). Similarly, if one assumes the latent dyanmic process in Section 5.2.2 is governed by a NN, then as in the case of B-PINNs, we have additional data models that condition on the neural network parameters as discussed in Section 4.2.
5.2 Process Models
We let the “process” part of the model be decomposed into three terms; an overall time-varying spatial mean, a latent dynamic process, and a residual (discrepancy) process, respectively as:
$\psi(\bm{u}_{t})=\bm{\mu}_{t}+\bm{\Phi}{\bm{u}}_{d,t}+\bm{\nu}_{t},$ (13)
where $\psi(\;)$ is a link function as in generalized linear mixed models (e.g., identity, logit, log, etc.) or an activation function as in neural modeling (e.g., sigmoid, tanh, ReLU, etc.), $\bm{u}_{d,t}$ is the latent dynamic process, and $\bm{\nu}_{t}$ is the discrepancy process. In this case, $\bm{\Phi}$ is a $n\times n_{d}$ matrix that can represent $n_{d}$ spatial basis functions, the identity (in which $n_{d}=n$), or some other linear mapping between the $\bm{u}_{d,t}$ dynamic process and the transformed process. Each of these terms can be specified in multiple ways as described below. Here, we assume $\bm{u}_{t}$ is an $n\times 1$ vector composed of the latent $u$-process at $n$ spatial locations (but as noted in Section 2.1, this can be a multivariate latent vector).
5.2.1 Mean
The mean process $\bm{\mu}_{t}$ can accommodate covariates or other “fixed effects”; e.g., $\bm{\mu}_{t}=\bm{X}_{t}\bm{\beta}_{t}$, where $\bm{X}_{t}$ is an $n\times q$ covariate matrix with potentially time-varying covariates, and $\bm{\beta}_{t}$ is a $q$-dimensional vector (that may be time varying). In this case, we would specify a prior distribution $[\bm{\beta}_{t}]$ (which depends on the problem). We use the term “covariates” quite loosely here, and note that $\bm{X}_{t}$ this term can accommodate seasonality, temporal trends, one hot encoding, exogenous covariate effects, etc. In some cases, one may not need this term in the model, or it might not depend on time (e.g., it might represent a simple constant offset).
5.2.2 Latent Dynamic Process
The latent dynamic component of the model ${\bm{u}}_{d,t}$ is the most important as it represents the underlying spatio-temporal dynamic mechanism driving the process. This is the main place where we impart mechanistic information (e.g., a PINN or a mechanistically-motivated spatio-temporal statistics model). Note the $n\times n_{d}$ matrix $\bm{\Phi}$ in (13) can represent a mapping matrix or, as is common in statistical DSTMs, a matrix of spatial basis functions (typically, low-rank). In that case, the $n_{d}$-dimensional dynamical process ${\bm{u}}_{d,t}$ operates in some basis space (e.g., see the overview in Cressie & Wikle,, 2011).
The dynamical process can be a PINN as described in Section 4.1 or B-PINN as in Section 4.2, in which case we write with a slight abuse of notation,
${\bm{u}}_{d,t}={\bm{u}}_{NN,t}(\bm{\theta}_{W}),$ (14)
where we are not explicitly writing out the initial and boundary conditions, but those can be incorporated in the PINN as described previously. In this case, the PDE constraint is incorporated into “observations” of the PDE residuals (8) as before (e.g., we can include a data model for these residuals and specify the observations to be zero unless there is a known PDE forcing term, or specify a prior on the time derivative of the PDE). This then accommodates the PDE dynamical structure and we note there are additional $\bm{\theta}_{m}$ parameters that come in the residual process from the PDE. As with the B-PINN framework, the residual variance term ($\sigma^{2}_{r}$) allows one to control the degree of determinism in the PDE representation. In full generality it should be emphasized that the latent dynamic process need not be deterministic. For example, as in Section 3.2.1 consider a mechanistically-motivated DSTM, which we write here as
${\bm{u}}_{d,t}={\cal M}({\bm{u}}_{d,t-1},B({\bm{u}}_{d,t};\bm{\theta}_{b});\bm{\theta}_{m})+\bm{\eta}_{t},$ (15)
where ${\cal M}(\;)$ represents a linear or nonlinear operator that depends on parameters $\bm{\theta}_{m}$ and boundary conditions $B({\bm{u}}_{d,t};\bm{\theta}_{b})$ (that can be learned), and the relationship is explicitly Markovian with a zero-mean Gaussian error process $\bm{\eta}_{t}$ with covariance matrix (potentially time-varying) $\bm{Q}_{t}(\bm{\theta}_{q})$ that depends on parameters $\bm{\theta}_{q}$ (e.g., for an overview see Cressie & Wikle,, 2011; Wikle et al.,, 2019). We also must specify explicitly an initial condition for the the process, say ${\bm{u}}_{d,0}\sim\text{Gau}(\bm{a},\bm{\Sigma}_{0})$ as is usual with spatio-temporal dynamic models.
An important difference between the PINN formulation in (14) and the mechanistically-motivated DSTM in (15) is that the NN model considers all times jointly (i.e., treating time as exchangeable), whereas the dynamic model is Markovian and explicitly considers time conditionally.
5.2.3 Discrepancy Process
The discrepancy process term $\bm{\nu}_{t}$ is important because it respects the fact that our latent dynamical model represented by the process ${\bm{u}}_{d,t}$ surely does not completely account for all of the latent spatio-temporal dependence in real-world data (i.e., “all models are wrong”). This term could be accommodated by another NN, say $\bm{\nu}_{t}=\bm{\delta}_{t}(\cal{S};\bm{\theta}_{\delta})$ as in Zou et al., (2024). Alternatively, it could be a time-varying stochastic process, e.g., $\bm{\nu}_{t}=f_{\nu}(\bm{\nu}_{t-1};\bm{\theta}_{\nu})$ such as the residual autoregresssive process to accommodate small scale dynamics in Wikle et al., (2001), or simply a time-independent spatial GP (e.g., $\bm{\nu}_{t}\sim\text{Gau}(\bm{0},\bm{\Sigma}))$. More generally, this term could be represented by any spatio-temporal random effect distribution, say $\bm{\nu}_{t}\sim[\bm{\theta}_{\nu}]$.
5.3 Parameter Models
Depending on the models selected in the data and process stages, one can have a variety of parameters that require prior distributions, e.g., $[\bm{\beta}_{t}|\bm{\theta}_{\beta}][\bm{\theta}_{w}][\bm{\theta}_{\nu}][\bm{\theta}_{\delta}],etc.$ In most cases, the parameters closer to the data and process should have informative priors, either because one has scientific knowledge (e.g., a diffusion parameter should change with space and/or time) or because they might be weakly identified and require empirical Bayes or meta-learning priors (e.g., as in the case of PINNs). Typically, we require regularization for these priors as there are likely to be quite a lot of them, or as mentioned in the context of physical-statistical models, we might have mechanistic information for processes (e.g., parameters must obey a turbulent scaling relationship as in Wikle et al., (2001)).
Deep BHMs as described here have multiple hyperparameters. These correspond to specifications of the parameters for the priors in the previous stages as well as model architecture (e.g., number of hidden units, levels, etc.) in the case of NN process models. Although in some cases these can be given vague proper prior distributions, they are typically fixed (i.e., point mass priors) and the sensitivity to the choice is examined. These choices are very problem specific.
5.4 Implementation Considerations
When implementing hybrid neural-statistical models such as GMID-DSTM with a NN dynamic process, multiple factors need to be considered. As with many mathematical and statistical models, model identifiability must be be addressed, specifically with regard to the NN and discrepancy term. For example, in (13), if $\bm{u}_{d,t}$ is parameterized using (14) (with a mechanistic component) and $\bm{\nu}_{t}$ is parameterized using a GP, there is potential for confounding between the two processes. There are multiple ways to mediate this issue. The most natural is to restrict the spatial (temporal, or spatio-temporal) range of the GP to operate on a different scale than the dynamics imposed on $\bm{u}_{d,t}$. This can be accomplished by an informative prior distribution on the argument(s) of the kernel function for the GP.
The next considerations are: how best to incorporate a NN within a statistical model, how to specify the architecture of the NN, and how to select prior distributions on the NN weights (and biases). As discussed in Section 4.2, a Bayesian NN (and by extension the B-PINN) is the most natural method to incorporate the NN. However, all the same challenges facing Bayesian NNs and B-PINNs are exhibited in the hybrid models. Regarding the NN architecture, this is a relatively unexplored area, though Toscano et al., (2025) reviews various neural network architectures that have been considered in the context of PINNs. Within the B-PINN literature, simple NNs, such as narrow 3-layer feed-forward NNs, have been used. This could be an artifact of the bias-variance trade-off, where wide and deep NNs are over-parameterized and have the potential to overfit. As such, empirical results suggest narrow and shallow feed-forward NNs result in the best performance, although this may not be the case for complex dynamics. Another potential reason for using relatively simple NNs in B-PINNs is they are more computationally tractable to fit, and we next discuss computational considerations.
The primary computational tool that has been used for posterior inference in Bayesian hierarchical models is MCMC, such as Gibbs sampling (and variants like Metropolis-within-Gibbs sampling) (Brooks et al.,, 2011). Modern gradient-based methods, such as Hamiltonian Monte Carlo (HMC; Neal et al.,, 2011) and extensions like the No-U-Turn Sampler (Hoffman & Gelman,, 2014), have become prevalent in general for Bayesian inference. Often it is the case that the computational cost of an expensive simulator makes MCMC prohibitively expensive, but an emulator or surrogate model can speed up the computations (Higdon et al.,, 2008; Hooten et al.,, 2011; Gopalan & Wikle,, 2022), especially in physical-statistical models. Moreover, physical-statistical models typically involve GPs, which can be computationally difficult due to cubic complexity of inversion and determinant calculations; however, there are ways to reduce this burden, such as the use of sparse precision matrices, reduced-rank basis representations, nearest neighbor approaches, and Vecchia approximations (e.g., see the review in Heaton et al.,, 2019).
In the B-PINN literature (Yang et al.,, 2021), the two primary computational tools are HMC, which has already been introduced, and variational inference (Blei et al.,, 2017). In the case of variational Bayes, the mean-field approximation (i.e., product of independent Gaussians) for posterior neural network weights might not be realistic in the sense that correlations (which are likely to exist in neural network parameters) are neglected, although variational GPs (Tran et al.,, 2015) can help alleviate this issue. Yang et al., (2021) found gradient-based MCMC methods result in the most optimal performance for B-PINNs and for our example presented below in Section 6, we proceed with this approach.
6 Simulation Experiment: GMID-DSTM
To demonstrate the GMID-DSTM framework presented in Section 5 with a B-PINN dynamic process, we consider a simulation experiment. In particular, we consider a mechanistic dynamic process described by a one-dimensional (in space) Burgers’ equation (Bateman,, 1915; Burgers,, 1948), which is a classic nonlinear PDE given by
$\frac{\partial\tilde{u}(s,t)}{\partial t}+\tilde{u}(s,t)\frac{\partial\tilde{u}(s,t)}{\partial s}-\lambda\frac{\partial^{2}\tilde{u}(s,t)}{\partial s^{2}}=0,$ (16)
where $\lambda$ is the diffusion coefficient. The Burgers’ equation is often used to illustrate dynamics because it is a very simple nonlinear advection-diffusion representation of a fluid dynamics problem that can model features like nonlinear wave propagation and shock formation while remaining relatively easy to analyze and solve. Here, we also include a fixed-effects mean term in the model and a Gaussian process discrepancy to reflect a situation where the true process is not just a known mechanistic model. We also add measurement error to the simulation process. The details of the simulation are given below.
Figure 1: Simulated process $u(s,t)$ (left) and components over one-dimensional space and time: Burgers’ equation $\tilde{u}(s,t)$ (second from left), mean process $\mu(s,t)$ (second from right), and discrepancy process $\nu(s,t)$ (right).
Figure 2: Left: Simulated $\bm{z}_{t}$, Right: Missing data in $\bm{z}_{t}$ highlighted in white.
  1. 1.
    Domain: We consider a one-dimensional equally spaced spatial grid with $n=51$ locations in the domain $s\in[-\pi,\pi]$ and $T=25$ equally distributed time points, $t\in[0,5]$.
  2. 2.
    Process Mean ($\mu(s,t)$): We consider two covariates in the mean process, $\bm{x}(s,t)=(x_{1}(s,t),x_{2}(s,t))^{\top}$, which are simulated from a zero-mean spatio-temporal GP with exponential covariance function $k_{x}((s,t),(s^{\prime},t^{\prime}))=0.1\cdot\exp\bigg{(}-\frac{\sqrt{(s-s^{\prime})^{2}+(t-t^{\prime})^{2}}}{\ell_{s}}\bigg{)}$ where $\ell_{s}$ is set to $1/20$ the maximum observed spatio-temporal distance across all location-time pairs. Specifically, we choose $\ell_{s}=8.03/20$ and let $\bm{\beta}=(0.3,-0.2)^{\top}$.
  3. 3.
    Latent Dynamic Process ($\tilde{u}(s,t)$): For each spatial and temporal location, we generated the PDE solution to Burgers’ equation (16) numerically (with a spectral solver), denoted by: $\tilde{u}(s,t)$. This numerical solution used initial condition $\tilde{u}(s,0)=\exp(-s^{2})$ and boundary conditions $\tilde{u}(-\pi,t)=\tilde{u}(\pi,t)=0$.
  4. 4.
    Discrepancy Process ($\nu(s,t)$): For each time point a spatial discrepancy field was drawn as an independent, zero-mean Gaussian process, i.e., $\nu(s,t)\sim\text{GP}({0},k(s,s^{\prime}))$, where $k(s,s^{\prime})=\sigma^{2}_{\nu}\exp\bigg{(}-\frac{(s-s^{\prime})^{2}}{2\ell_{\nu}^{2}}\bigg{)}$ is a squared-exponential covariance function. Specifically, we select $\sigma^{2}_{\nu}=0.05$, and $\ell_{\nu}=0.15$.
  5. 5.
    Simulated Process ($u(s,t)$): The simulated process at each spatial and temporal coordinate is then given as the sum of the mean, latent dynamic process, and discrepancy: $u(s,t):=\bm{x}(s,t)^{\top}\bm{\beta}+\tilde{u}(s,t)+\nu(s,t)$.
  6. 6.
    Simulated Observations ($z(s,t)$): We assumed that 50% of the space coordinates are missing (sampled at random spatially and fixed across all time points - which is generally a realistic scenario for simulating environmental processes compared to data being missing completely at random in space and time). We then added mean zero independent, identically distributed Gaussian measurement error to the process values, $u(s,t)$ for the randomly selected spatial and temporal coordinates with measurement error variance $\sigma^{2}_{d}=0.2^{2}$, to get observations $z(s,t)$.
Figure 1 shows the simulated space-time fields of the true process, latent dynamic process, mean, and discrepancy fields with the simulated data and data with missing observations given in Figure 2.
We consider the following BHM to fit these simulated data. First, the data and process models are given by:
[…]
The latent dynamic and discrepancy processes are then given by:
[…]
$\quad\frac{\partial{u}_{NN}(s,t)}{\partial t}\sim\text{Gau}(u_{NN}(s,t)\frac{\partial{u}_{NN}(s,t)}{\partial s}-\lambda\frac{\partial^{2}{u}_{NN}(s,t)}{\partial s^{2}},\sigma^{2}_{r})$
[…]
$\quad k(s,s^{\prime})=\sigma^{2}_{\nu}\exp\bigg{(}-\frac{(s-s^{\prime})^{2}}{2\ell_{\nu}^{2}}\bigg{)},$
In this case, $\bm{H}_{t}$ is an incidence matrix that accommodates missing observations (e.g., see the discussion in Wikle et al.,, 2019). In addition, the following prior distributions are specified:
$\sigma^{2}_{d}$ $\sim\text{Cauchy}(\mu_{d},\gamma_{d})\cdot\text{I}(\ell_{d}<\sigma^{2}_{d}<u_{d}),\qquad$ $\bm{\beta}\sim\text{Gau}(\bm{\mu}_{\bm{\beta}},c_{\bm{\beta}}\bm{I})$
$\lambda$ $\sim\text{Gau}(\mu_{\lambda},\sigma^{2}_{\lambda})\cdot\text{I}(0\leq\lambda<\infty),\qquad$ $\sigma^{2}_{\nu}\sim\text{Cauchy}(\mu_{\nu},\gamma_{\nu})\cdot\text{I}(\ell_{\nu}<\sigma^{2}_{\nu}<u_{\nu})$
$\ell_{\nu}$ $\sim\text{Gau}(\mu_{\ell},\sigma^{2}_{\ell})\cdot\text{I}(\ell_{\ell}\leq\ \ell_{\nu}<u_{\ell}),\qquad$ $\bm{\theta}_{W}\sim\text{Gau}(\bm{\mu}_{W},c_{W}\bm{I}),$
where relatively informative priors are assigned for $\sigma^{2}_{d}$, $\sigma^{2}_{\nu}$, and $\ell_{\nu}$ with $\mu_{d}=\mu_{\nu}=\mu_{\ell}=0,\gamma_{d}=\gamma_{\nu}=\sigma^{2}_{\ell}=1$, and truncation bounds $\ell_{d}=0.1,u_{d}=0.3,\ell_{\nu}=0.02,u_{\nu}=0.1,\ell_{\ell}=0.05,$ and $u_{\ell}=0.2$. This is to address the potential identifiability issue between the measurement error $\sigma^{2}_{d}$ and the discrepancy scale parameter, as well as the ability to recover Gaussian process covariance parameters in a fixed domain setting (an issue which has been frequently noted in the spatial statistics literature as for instance in Gelfand et al., (2010)). In contrast, we assign relatively non-informative priors to $\bm{\beta}$, $\lambda$, the neural network weight and bias parameters $\bm{\theta}_{W}$, with $\bm{\mu}_{\bm{\beta}}=\bm{\mu}_{W}=\bm{0}$, $c_{W}=\sigma^{2}_{\lambda}=1$, $\mu_{\lambda}=0$, and $c_{\bm{\beta}}=10$. The neural network for ${\bm{u}}_{NN,t}$ consists of $L=3$ hidden layers, each with 16 hidden units.
Figure 3: Violin plots of the BHM model parameter posterior distributions, with the true value indicated by a dashed red line. From left to right, the plots correspond to $\sigma^{2}_{\nu}$, $\ell_{\nu}$, $\lambda$, $\sigma^{2}_{d}$, $\beta_{1}$, and $\beta_{2}$.
Figure 4: Left: true $\tilde{\bm{u}}_{t}$; Second from the left: Predicted $\tilde{\bm{u}}_{t}$ from $u_{NN}(s,t)$; Second from the right: Posterior means (dashed line) and corresponding 95% credible intervals (shaded region) for predicted values of a marginal spatial slice of $\tilde{\bm{u}}_{t}$ at times $t=0$, $2.5$, and $5$ compared to the true process (solid line); Right: Posterior means (dashed line) and corresponding 95% credible intervals (shaded region) for a marginal time slice of the predicted values of $\tilde{\bm{u}}_{t}$ at spatial locations $s=-3.14$, $-1.13$, and $1.01$ compared to the true process (solid line). In the right two plots, the dots correspond to spatial locations or time points when data are observed.
Figure 5: First Row, First Column: True values of the spatio-temporal discrepancy term $\bm{\nu}_{t}$; Second Row, First Column: Corresponding posterior prediction of $\bm{\nu}_{t}$; First Row, Columns 2-4: Spatial slice of posterior mean (dashed line) and 95% credible intervals (shaded region) for times $t=0$, $2$, and $t=5$ compared to the true process (solid line); Second Row, Columns 2-4: Time slice of posterior mean (dashed line) and 95% credible intervals (shaded region) for spatial locations $s=-3$, $-1$, and 1 compared to the true process (solid line). For all plots in columns 2-4, the dots correspond to spatial locations or time points when data is observed.
As discussed in Section 4.2, the specification of the B-PINN model in this BHM requires PDE residuals, boundary condition data, and initial condition data. Specifically, we can write the following conditional data distributions that have the effect of updating the prior distributions for weights $\bm{\theta}_{W}$.
[…]
$[D_{bc+}|\sigma^{2}_{bc},\bm{\theta}_{W}]$ $=\text{Gau}(0;u_{NN}(\pi,t;\bm{\theta}_{W}),\sigma^{2}_{bc}),\quad(\text{Right boundary condition})$
$[D_{bc-}|\sigma^{2}_{bc},\bm{\theta}_{W}]$ $=\text{Gau}(0;u_{NN}(-\pi,t;\bm{\theta}_{W}),\sigma^{2}_{bc}),\quad(\text{Left boundary condition})$
$[D_{ic}|\sigma^{2}_{ic},\bm{\theta}_{W}]$ $=\text{Gau}(\exp(-s^{2});u_{NN}(s,0;\bm{\theta}_{W}),\sigma^{2}_{ic}),\quad(\text{Initial condition}),$
where $\sigma^{2}_{r}$, $\sigma^{2}_{bc}$, and $\sigma^{2}_{ic}$ govern the extent to which the neural network $u_{NN}$ is constrained to satisfy Burgers’ equation along with the initial conditions and boundary conditions. Assuming reasonably strong prior information for the PDE, boundary conditions, and initial conditions, we set $\sigma^{2}_{r}=0.05^{2}$ and $\sigma^{2}_{bc}=\sigma^{2}_{ic}=0.01^{2}$. We emphasize that assuming the residual data are zero (e.g., $D_{r}=0$) further constrains the prior on $\bm{\theta}_{W}$ and these residuals are not actual measurements of the process.
After specifying the prior distributions and the neural network architecture, we draw 2,000 posterior samples using HMC with the No-U-Turn Sampler (Hoffman & Gelman,, 2014), discarding the first 1,000 as burn-in and using a tree depth of 10. This is implemented using the “NumPyro" (Phan et al.,, 2019) and “JAX" (Bradbury et al.,, 2018) packages in Python and takes approximately 90 minutes on an Apple MacBook M1 Pro with 16GB RAM.
Figure 6: Top Left: True simulated $\bm{u}_{t}$; Top Right: Predicted posterior mean $\bm{u}_{t}$; Bottom Left: Posterior standard deviation, with white rectangles indicating locations where data are observed; Bottom Right: Difference between the true $\bm{u}_{t}$ and the posterior mean.
6.1 Results
We first evaluate whether the posterior samples generated from the No-U-Turn sampler produce accurate estimates of the true simulation parameters. Figure 3 presents violin plots for the parameters, excluding the neural network weights and biases $\bm{\theta}_{W}$, which are not the primary focus. Although some parameters show slight estimation bias, their credible intervals include the true values and are indicative of Bayesian learning.
Figures 4 and 5 compare the true values and predictions of ${\bm{u}}_{d,t}$ and $\bm{\nu}_{t}$, respectively. As shown in Figure 4, the predicted ${\bm{u}}_{d,t}$ space-time field and spatial and temporal marginal plots align well with the true values, despite the substantial amount of missing data. This is one of the primary advantages of incorporating mechanistic information into the model. In contrast, Figure 5 demonstrates that accurately predicting the discrepancy term, $\bm{\nu}_{t}$, is more challenging and subject to greater uncertainty, although the posterior mean (dashed line in the spatial and temporal slice plots) still reasonably matches the true process (solid line). Combining posterior samples from each term, Figure 6 shows the posterior mean of the overall process $\bm{u}_{t}$. While there are slight deviations from the true $\bm{u}_{t}$, the posterior mean successfully captures the underlying spatio-temporal pattern, demonstrating the flexibility of the model.
7 Discussion
The incorporation of mechanistic information in data-driven models has been of interest to scientists for decades. There are numerous paradigms for such data-driven approaches, ranging from variational data assimilation, online filtering, physical-statistical models and probabilistic numerical methods. Recently, there has been a resurgence of interest in these approaches because of the success of using deep neural networks with physical constraints. Here, we review these different approaches and show some connections, followed by presenting a general BHM framework that can accommodate each of these paradigms. We illustrate this with a simulation of a process that is informed by the classic Burgers’ equation nonlinear PDE, and demonstrate we can recover important components of the model (fixed effect mean parameters, the Burgers’ diffusion parameter, and a discrepancy process) given uncertain and missing data.
One novel contribution of this work has been the introduction of a mechanistically-informed BHM (GMID-DSTM) that considers a B-PINN for Burgers’ equation (Section 6). We expect that the same sort of approach can generalize to different processes motivated by alternate PDEs. In other cases, it may be preferred to embed a dynamical process that has spatio-temporal covariance, in contrast to when dynamics are encoded with a neural network. For example in Wikle et al., (2001), 2-D spatio-temporal wind velocity follows a vector-autoregressive process that is motivated by an underlying set of (linear) PDEs. However, for general nonlinear PDEs there may not be an autoregressive or general quadratic nonlinear (Wikle & Hooten,, 2010) representation or approximation of the latent dynamics. In such cases a mechanistic-BHM using a PINN approach may be preferred since automatic differentiation can be used for general, nonlinear PDEs. This could also be an advantage of using a model such as the one in Section 6 over PNM methods that lack a neural network representation, in which case it is not obvious that automatic differentiation would apply.
From our exploration of these methods, in particular, the incorporation of a B-PINN within a more general BHM, there are several challenges with this approach that suggest additional research. Of primary concern is the challenge of computational efficiency when scaling this to a realistic environmental application. Although automatic differentiation is quite helpful for implementing PDE models, it can be slow, and there are well-known challenges to Bayesian learning of deep neural network parameters. In addition, real-world processes are often described by more complex dynamics than can be parameterized by a PDE (all models are wrong), and modeling the discrepancy becomes crucial. Our simulation studies have shown there can be significant identifiability issues with a flexible discrepancy process and the underlying mechanistic process without rough a priori guidance as to at least the space and time scales corresponding to each. It is also the case the data volume and orientation can make a difference in accurately estimating the various model components. Lastly, as with all deep neural implementations, there is typically no a priori guidance to inform the model architecture. These are very active areas of research and we are optimistic that these challenges can be met and these methods can be scaled to many realistic spatio-temporal problems.
Acknowledgments
JSN was supported by the Director, Office of Science, Office of Biological and Environmental Research of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 and by the Regional and Global Model Analysis Program area within the Earth and Environmental Systems Modeling Program as part of the Calibrated and Systematic Characterization, Attribution, and Detection of Extremes (CASCADE) project.
GG was supported by the Verification and Validation subprogram of the Advanced Simulation and Computing Program at the Los Alamos National Laboratory.